Note:      1. Answer all questions in space provided.

     2. Show your work to receive full marks.

     3. Be sure to upload your completed work as Word Document;

          no other format will be accepted including .pages.

 

Question 1:

How much interest will you pay on a debt of $7,500 if you are paying the loan off in nine months. The interest rate is 4.375%. ______________

 

 

<b><span style="font-family:&quot;Verdana&quot;,sans-serif">Question 2:<o:p></o:p></span></b>

ABC Corp. invested $13,500 in a 270-day term deposit. What is the maturity value if the rate of interest is 3.65%? How much interest will the term deposit earn? ________________

<b><span style="font-family:&quot;Verdana&quot;,sans-serif"> </span></b>

 

 

Question 3:

What amount of money that, deposited in an account on April 1, 2021, will grow to $657.58 by September 10, 2021, at 4.75% p.a.?     __________________

 

 

 

Question 4:

When Mark borrowed $2,300, he agreed to repay the loan in two equal payments, to be made 90 days and 135 days from the day the money was borrowed. If interest on the loan is 9.25%, what is the size of the equal payments if a focal date of today is used? _______________<b> <o:p></o:p></b>

 

 

 

 

Question 5:

What is the price of a $25,000, 91-day Province of British Columbia Treasury bill on its issue date if the current market rate of return is 7.672% p.a.? ______________

 

 

 

 

The End



1
Expert's answer
2021-02-25T01:26:29-0500

solution 1. Debt amount "= \$ 7500"

Rate of interest "(r)=4.375\%"

Hence, the required interest "=\frac{7500\times 4.375\times 9}{12\times 100 }"

"=\$ 246.0"

Solution 2.

"SI=\frac{13500\times 3.67\times 270}{100\times 365}=366.5\$"

solution 3:

"SI=\frac{PRT}{100}"

Solution 2.

Principal amount "(P)=\frac{12\times 100\times 657.58 \times 365 }{4.75\times 40}"

=1515894.94$

Solution 4

Let one payment be x, so

"2300=[x-x\times 0.0925\times \frac{135}{365}]+ [x-x\times 0.0925\times \frac{90}{365}]"

"2300=1.942x"

"x=1184.34\$"


Solution 5

Interest "=\frac{25000\times 91\times 7.672}{100\times 365}"

"=\$478"

Hence, final price of the product "=\$25000+\$478"

"=\$ 25478"